OpenStudy (anonymous):

Adding, Subtracting, Multiplying, and Dividing Functions. Can someone help me with this subject

OpenStudy (anonymous):

I might be able to but depends on the question though

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Two neighboring towns, Nash Town and Bay Town, were surveyed, and their populations were found to be x1/3 +2x1/5 and 3x1/3 - x1/5 (in thousands), respectively, where x is the number of years since the towns were founded. (Assume that both towns were founded at the same time.)

OpenStudy (anonymous):

Question 1 Find the function that gives the total population of both towns.

OpenStudy (anonymous):

@Anna-Rose , @Arn.01 , @nincompoop, @BioHazard9064

OpenStudy (anonymous):

@Barrelracing, @freegirl11112, @iloveyou;*, @jaredhm29, @Nnesha, @Tashiya_Pandas

OpenStudy (anonymous):

sorry i suck at math :/

OpenStudy (anonymous):

its alright

OpenStudy (biohazard9064):

merda

OpenStudy (biohazard9064):

Due città vicine , Nash città e sulla baia di Città , sono state intervistate , e le loro popolazioni sono risultati x1 / 3 + ​​2x1 / 5 e 3x1 / 3 - x1 / 5 (in migliaia ), rispettivamente, dove x è il numero di anni da quando le città sono state fondate . ( Si supponga che entrambe le città sono state fondate allo stesso tempo . )

OpenStudy (anonymous):

did youjust translate that

OpenStudy (anonymous):

por que

OpenStudy (biohazard9064):

perché

OpenStudy (biohazard9064):

ok

OpenStudy (biohazard9064):

Did you try google

OpenStudy (biohazard9064):

nvm

OpenStudy (biohazard9064):

@LynFran Puoi Aiutare

OpenStudy (biohazard9064):

I dont know

OpenStudy (biohazard9064):

scusate - sorry

OpenStudy (biohazard9064):

Two neighboring towns, Nash Town and Bay Town, were surveyed, and their populations were found to be x1/3 +2x1/5 and 3x1/3 - x1/5 (in thousands), respectively, where x is the number of years since the towns were founded. (Assume that both towns were founded at the same time.)

OpenStudy (biohazard9064):

ok so lets break it down a little what do we need to know and what we dont

OpenStudy (anonymous):

sorry @BioHazard9064 cant help

OpenStudy (biohazard9064):

its ok

OpenStudy (biohazard9064):

So Nash Town would be x1/3 +2x1/5 and (in thousands) Bay Town 3x1/3 - x1/5 And x = years

OpenStudy (biohazard9064):

after that i dont know

OpenStudy (jacob902):

Example Problem f(x) = 2x g(x) = 3x+1 What is f(3) + g(4) ? Step 1: evaluate the individual functions f(3) = 2×3 = 6 g(4) = 3×4 +1 = 13 Step 2: Add the values that your functions produced f(3) + g(4) = 6 + 13 = 19

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@nincompoop

OpenStudy (dan815):

okay so

OpenStudy (dan815):

i think the main operation to understand here is addition the rest stem from it

OpenStudy (dan815):

lets start with a simple example okay now u know the way they teach u about adding 2 numbers 111 +234 ------ 345 They teach you to add by column, and carry when u pass 10

OpenStudy (dan815):

are you with me so far

OpenStudy (anonymous):

yes

OpenStudy (dan815):

now why column by column its because u can also think about it like adding 'like' terms

OpenStudy (dan815):

you are adding the 1s units, the 10s units and 100s units together

OpenStudy (anonymous):

i need help putting it in an equation then i can solve it

OpenStudy (dan815):

in otherwords \[111+234= (100+10+1) + (200+30+4)\\ =(1*10^2+1*10^1+1*10^0) +(2*10^2+3*10^1+4*10^0)\] now you add these like terms like so \[1*10^2+2*10^2 + 1*10^1 + 3*10^1+1*10^0+4*10^0\\ =(1+2)*10^2+(1+3)*10^1+(1+4)*10^0\] this is very analogous to what polynomial addition instead of 10, wwe have x there

OpenStudy (anonymous):

oh thx

OpenStudy (anonymous):

oh thx